OpenStudy (anonymous):

Can you give an example of a periodic function f(g(x)) - Where f(x) and g(x) are each non periodic?

OpenStudy (zarkon):

\[f(x)=\sin(x^3)\] \[g(x)=\sqrt[3]{x}\]

OpenStudy (anonymous):

Haha. Nice really. :)